OpenStudy (anonymous):

The defining equations for the function f is f(x) = 3x^2 + 1. The defining equation for the function g is g(x) = 2x-3. Then the defining equation for f \circle g (x) is:

OpenStudy (anonymous):

10x^2-5x-5 12x^2-36x+28 6x^2 + 28 5x^3 - 12x^2 + 9x - 1

OpenStudy (anonymous):

@phi @binks

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

\[\Large f \circ g\] is the same as f(g(x))

jimthompson5910 (jim_thompson5910):

ie \[\Large (f \circ g)(x) = f(g(x))\]

jimthompson5910 (jim_thompson5910):

\[\Large f(x) = 3x^2 + 1\] \[\Large f(g(x)) = 3(g(x))^2 + 1\] \[\Large f(g(x)) = 3(2x-3)^2 + 1\] \[\Large f(g(x)) = 3(2x-3)(2x-3) + 1\] \[\Large f(g(x)) = 3(4x^2-12x+9) + 1\] \[\Large f(g(x)) = 12x^2-36x+27 + 1\] \[\Large f(g(x)) = 12x^2-36x+28\] \[\Large (f \circ g)(x) = 12x^2-36x+28\]